# LOGCAN++: LOCAL-GLOBAL CLASS-AWARE NETWORK FOR SEMANTIC SEGMENTATION OF REMOTE SENSING IMAGES

### v1

@@ V-D3 Analysis of the patch number
16 86.20 91.90 72.01 80.78 79.43 35.18 74.25
V-D4 Analysis of the head number
We finally investigate the effect caused by different numbers of heads on context modeling in the LCA module, where we set the number of heads as 1, 2, 4, 8 and 16 for the experiment. As shown in Table VII, the number of heads increases, the performance of the model tends to increase and then decrease, and the phenomenon is similar to the multi-head self-attention in Transformer [33]. This proves that by introducing the multi-head mechanism, LOGCAN++ is able to capture more complex pixel class relationships, thus improving the model performance.
@@ VI Conclusion
In this paper, we begin with a detailed analysis of the characteristics of remotely sensed images (i.e., complex background, scale and orientation variations, and high intra-class variance). Against this background, we propose the highly customized LOGCAN++ to cope with the segmentation challenges posed by the above characteristics. Specifically, LOGCAN++ is based on a local global class-aware strategy to reduce background noise interference and mitigate the intra-class variance problem. In particular, LOGCAN++ can tolerate scale and orientation variations in remote sensing images well due to the introduction of affine transformations in the process of extracting local class centers. Experimental results on three remote sensing semantic segmentation datasets show that LOGCAN++ significantly outperforms other state-of-the-art methods and possesses a better balance between accuracy and efficiency. In our future work, we will further explore it in conjunction with existing semantic segmentation macromodels such as SAM in order to fully utilize the potential of LOGCAN++ in the field of remote sensing image segmentation.

### v3

@@ V-D4 Analysis of the patch number
[…]
V-D5 Analysis of the head number
We investigate the effect caused by different numbers of heads on context modeling in the LCA module, where we set the number of heads as 1, 2, 4, 8 and 16 for the experiment. As shown in Table VIII, the number of heads increases, the performance of the model tends to increase and then decrease, and the phenomenon is similar to the multi-head self-attention in Transformer [36]. This proves that by introducing the multi-head mechanism, LOGCAN++ is able to capture more complex pixel class relationships, thus improving the model performance.
TABLE IX: Ablation Study on the loss function coefficients on the Vaihingen Dataset. The highest value is bolded.
Coefficient of $\mathcal{L}_{ce}^{aux}$ Impervious surface Building low vegetation Tree Car Clutter mIoU
[…]
0.8 85.87 91.85 71.44 80.35 80.14 38.63 74.72
[…]
V-D6 Analysis of the loss function coefficients
We finally explore the impact of different loss coefficients on model performance. As shown in Table IX, when we set the weight of auxiliary loss to 0.8, the proposed method achieves the the highest segmentation performance. Additionally, we observe that setting this value too low or too high slightly degrades the model’s performance.
@@ VI Conclusion
In this paper, we begin with a detailed analysis of the characteristics of remote sensing images (i.e., complex background, scale and orientation variations, and high intra-class variance). Against this background, we propose the highly customized LOGCAN++ to cope with the segmentation challenges posed by the above characteristics. Specifically, LOGCAN++ is based on a local global class-aware strategy to reduce background noise interference and mitigate the intra-class variance problem. In particular, LOGCAN++ can tolerate scale and orientation variations in remote sensing images well due to the introduction of affine transformations in the process of extracting local class centers. Experimental results on three remote sensing semantic segmentation datasets show that LOGCAN++ significantly outperforms other state-of-the-art methods and possesses a better balance between accuracy and efficiency. In our future work, we will further explore it in conjunction with existing semantic segmentation macromodels such as SAM in order to fully utilize the potential of LOGCAN++ in the field of remote sensing image segmentation.